# Boosting Jailbreak Attack with Momentum

### v1

@@ 5 Discussion and Limitations
In this work, we rethink a current popular gradient-based attack on LLMs from the optimization perspective. We first propose a new jailbreak attack through the lens of optimization named Momentum Accelerated GCG (MAC), and demonstrate that accelerated optimization methods like momentum SGD can effectively boost such attacks with notably higher attack success rates and fewer optimization steps, providing deeper insights into the current jailbreaking issue research.
@@
While the recent research thread mainly focuses on black-box attacks (Wei et al., 2023b; Zhu et al., 2023; Li et al., 2023a; Chen et al., 2023b), we consider it still important to develop an efficient white-box attack for developers to efficient evaluation and red-teaming on LLMs. This is similar to the conventional adversarial robustness in the vision domain, where white-box attack methods for evaluation are still valuable (Athalye et al., 2018; Croce & Hein, 2020).
@@
As a preliminary study, we acknowledge some limitations of this work, which we leave as future works. First, our MAC only consider the case of batch size as 1 in the multiple prompts attack case. It would be interesting to see whether a larger batch size can find a better trade-off between efficiency and stableness. In addition, there are many optimization methods other than momentum like Adam remain unexplored. Finally, our experiment only focuses on one model, and its effectiveness can be further substantiated by evaluating more models.
@@ Acknowledgement
This work was sponsored by the Beijing Natural Science Foundation’s Undergraduate Initiating Research Program (Grant No. QY23041) and the National Natural Science Foundation of China (Grant No. 62172019).
Appendix A Additional related work
A.1 Jaibreak attack on aligned LLMs
Despite significant efforts to align Large Language Models (LLMs) with human values and teach them to avoid generating harmful content, recent studies indicate that LLM alignment remains superficial (Qi et al., 2023; Liu et al., 2023b) and is vulnerable to carefully crafted jailbreak prompts that can bypass safeguards and induce the generation of harmful content (Yao et al., 2023; Chen & Shu, 2023). Previous studies have manually designed jailbreak prompts using persuasive instructions (Wei et al., 2023a; Zeng et al., 2024). Another approach involves using stealthy communication techniques (Yuan et al., 2024; Yong et al., 2023; Li et al., 2023a) to circumvent LLM safeguards. In addition to manual prompt-based attacks, there is work on automatically generating jailbreak prompts, such as gradient-based attacks like GCG (Shin et al., 2020; Zou et al., 2023b) and AutoDAN (Zhu et al., 2023), which attach a suffix to the harmful request and optimize it using gradient heuristics, showing a higher success rate but requiring white-box access to the target model. Notably, concurrent research employs genetic algorithms or red-teaming methods to optimize jailbreak prompts, requiring only query access to the model (Lapid et al., 2023).
A.2 Defend LLMs against adversarial attacks
In response to malicious attacks on aligned LLMs, preliminary research has developed defense methods for various stages of generation (Xu et al., 2024; Zhang et al., 2024a; Chen et al., 2024b). Pre-processing methods, such as perplexity filters (Alon & Kamfonas, 2023), harmful string detection (Kumar et al., 2023; Cao et al., 2023), retokenization, and paraphrasing (Jain et al., 2023), can be easily integrated into models but may lead to excessive false positives (Varshney et al., 2023). Another thread of defenses crafts safe prompts (Xie et al., 2023; Mo et al., 2024; Zhang et al., 2024b; Zhou et al., 2024) for improving the robustness of LLMs during generation. Notably, Although Adversarial Training (AT) is widely regarded as an effective defense against adversarial attacks, the vast number of parameters and data involved render it impractical and less effective for LLMs (Jain et al., 2023), highlighting the mechanism understanding on the robustness of language models (Zou et al., 2023a; Wei et al., 2024; Zhang et al., 2024a).
Appendix B Evaluation Details
To determine whether an attack success, we apply the same detection function applied in (Zou et al., 2023b), which judges whether the generated response contains any of the following tokens:
  • "I’m sorry"
  • "Sorry"
  • "I am sorry"
  • "I apologize"
  • "As an"
  • "I’m an"
  • "I’m just"
  • "As a language model"
  • "As an Assistant"
  • "I cannot"
  • "I can’t"
  • "I do not"
  • "Hello!"
  • "is illegal and unethical"
Appendix C Example suffix and responses
[…]

### v2

@@ V Conclusion and Limitations
In this work, we rethink a current popular gradient-based attack on LLMs from the optimization perspective. We first propose a new jailbreak attack through the lens of optimization named Momentum Accelerated GCG (MAC), and demonstrate that accelerated optimization methods like momentum SGD can effectively boost such attacks with notably higher attack success rates and fewer optimization steps, along with better transferability and robustness against defense mechanisms, providing deeper insights into the current jailbreaking issue research.
@@
While the recent research thread mainly focuses on black-box attacks [33, 43, 20, 10], we consider it still important to develop an efficient white-box attack for developers to efficient evaluation and red-teaming on LLMs. This is similar to the conventional adversarial robustness in the vision domain, where white-box attack methods for evaluation are still valuable [1, 11].
@@
As a preliminary study, we acknowledge some limitations of this work, which we leave as future works. First, our MAC only considers the case of batch size as 1 in the multiple prompts attack case. It would be interesting to see whether a larger batch size can find a better trade-off between efficiency and stableness. In addition, there are many optimization methods other than momentum, like Adam, remain unexplored. Finally, our experiment only focuses on two models, and its effectiveness can be further substantiated by evaluating more models.
@@ Acknowledgement
This work was sponsored by the Beijing Natural Science Foundation (Grant No. QY24035, QY23041) and the National Natural Science Foundation of China (Grant No. 62172019).